### v1

# Policy Space Response Oracles: A Survey

Abstract
In game theory, a game refers to a model of interaction among rational decision-makers or players, making choices with the goal of achieving their individual objectives. Understanding their behavior in games is often referred to as game reasoning. This survey provides a comprehensive overview of a fast-developing game-reasoning framework for large games, known as Policy Space Response Oracles (PSRO). We first motivate PSRO, provide historical context, and position PSRO within game-reasoning approaches. We then focus on the strategy exploration problem for PSRO, the challenge of assembling an effective strategy portfolio for modeling the underlying game with minimum computational cost. We also survey current research directions for enhancing the efficiency of PSRO, and explore the applications of PSRO across various domains. We conclude by discussing open questions and future research.
1 Introduction
In recent decades, the exploration of multiagent systems has been a central focus in Artificial Intelligence (AI) research. A multiagent system, often referred to as a game, comprises multiple decision-making agents that interact within a shared environment. To understand the strategic behavior among these agents – where the optimal behavior of one agent depends on the behavior of others – game theory provides a mathematical framework that defines behavioral stability through solution concepts like the Nash equilibrium (NE). To identify such solutions, various equilibrium computation approaches have been developed von Stengel (2002); Savani and Turocy (2023), either to enumerate all equilibria (see the work by Avis et al. Avis et al. (2010) for bimatrix games), or to find a single sample equilibrium (e.g., with the Lemke-Howson algorithm for a bimatrix game, or Linear Programming for a zero-sum matrix game von Stengel (2002)).
As the size of the game (i.e., the number of players and strategies) grows, the computational feasibility of enumeration diminishes, and one tends to focus on finding a sample equilibrium. Unlike for general-sum games, for zero-sum games, a sample equilibrium already provides valuable insights into effective strategic play, for example, by revealing the game’s (unique) value, which can be achieved with a strong strategy irrespective of the strategy of the other player. However, even in zero-sum settings, many games that arise from practical applications are too large to solve for a sample equilibrium even with polynomial-time methods for Linear Programming – these games can be extremely large that they are infeasible to be explicitly represented as a payoff matrix in practice. It is these huge games that are our primary focus in this survey.
As an alternative to traditional equilibrium computation methods, to reason about such huge games, a wide range of learning methods have been applied (with one of the most prominent approaches being multiagent Reinforcement Learning (RL) Albrecht et al. (2024)). While learning approaches have significantly contributed to the development of intelligent agents, they face many inherent challenges in games. For example, independent learning across agents can render the environment non-stationary, which is a challenge for convergence as each individual learner faces a potentially moving target. Another challenge is non-transitivity of a game, where there is not a clear notion of “better” strategy for an agent, and thus effective learning requires the learning scheme to maintain a population of strategies for each agent.
Against this backdrop, the Policy Space Response Oracles (PSRO) framework Lanctot et al. (2017) emerged as a natural blend that combines traditional game-theoretic equilibrium computation with learning. In PSRO, a key concept is an empirical game model, which acts as an approximation of the underlying full game. Game models are induced from simulations run over combinations of a particular set of strategies. This set of strategies is typically much smaller than the full game, and thus game models are feasible to analyze with traditional equilibrium computation methods. PSRO alternates between the analysis of the current game model, defining a new learning target, and game model refinement by including the new strategies generated via learning.
As a general solver for large-scale games, PSRO has been successfully applied to a wide range of game types and diverse application domains, from mechanism design Zhang et al. (2023) to robust RL Liang et al. (2023). Numerous PSRO variants have been developed, each tailored to leverage the specific characteristics of the underlying games. As a notable example of its success, algorithms inspired by PSRO have reached state-of-the-art performance in large-scale games such as Barrage Stratego McAleer et al. (2020), and in StarCraft Vinyals et al. (2019), where they have convincingly outperformed human experts and prior AI systems.
While PSRO and its variants have been covered to some extent in existing multiagent learning surveys (e.g., Yang and Wang (2020); Long et al. (2023)), a dedicated survey on PSRO like this one has been lacking. In this survey, we reflect on the historical development of PSRO, which arose from different research communities, and we place PSRO within the space of game reasoning approaches. We then present the latest developments on PSRO, highlighting both current and future research directions.
Figure 1: Illustration of the DO and PSRO frameworks respectively. The PSRO framework generalizes the DO framework by introducing MSSs, enabling best-response targets other than NE. Besides, PSRO accommodates various ROs and (approximate) best-response oracles.
2 The PSRO Framework
The PSRO framework incorporates a synthesis of ideas originating from distinct research communities. Within the planning community, McMahan et al. McMahan et al. (2003) laid the groundwork by formulating robust planning in Markov Decision Processes as a zero-sum game characterized by a vast strategy space. Drawing inspiration from Bender’s decomposition in optimization Benders (1962), they introduced the Double Oracle (DO) algorithm. DO is an iterative algorithm for solving games with a finite number of strategies. DO maintains a restricted version of the full game and iteratively expands the restricted game by adding best responses to the current equilibrium. When DO terminates, no player can deviate unilaterally to gain extra payoff and therefore the equilibrium in the current restricted game is an NE of the full game. In finite games, DO is guaranteed to converge to NE, though the restricted game could include all strategies of the full game in the worst case.
Concurrently, similar methods were being developed from a different perspective in the co-evolution research community. Co-evolutionary methods evolve multiple populations of (different) species in parallel. In this setting, there is no explicit fitness function, but the fitness of a population member depends on how well it interacts with members of other populations. Traditional challenges for co-evolution include the presence of intransitive cycles in traits, and the related problem of forgetting a trait that seems not useful at one point in the evolutionary process but later becomes useful again. The community explored memory mechanisms and game formulations where the co-evolutionary process was setup to discover NE as mixtures of traits Angeline et al. (1993); Popovici et al. (2012). To overcome the mentioned obstacles of intransitive cycles and forgetting traits, various archival structures were devised to facilitate monotonic convergence towards a range of game-theoretic solution concepts. For example, the “Nash memory” for symmetric games Ficici and Pollack (2003) identifies equilibrium strategies within a discovered restricted game. This was extended to asymmetric games via the Parallel Nash Memory Oliehoek et al. (2006), broadening the standard DO framework to accommodate alternative oracle types beyond the best response.
The PSRO framework Lanctot et al. (2017) was introduced in a paper titled “A Unified Game-Theoretic Approach to Multiagent Reinforcement Learning”, bringing many of the mentioned ideas from the planning, multiagent RL, and co-evolution communities together in a unified framework, which also covers classical game-theoretic learning dynamics such as Fictitious Play Brown (1951). Technically, PSRO generalizes DO by introducing the concept of meta-strategy solver (MSS), which extracts a profile from the current restricted game as the next best-response target (see Figure 1). This enables the best-response target to extend beyond NE, transforming PSRO into a versatile framework capable of generalizing various classical and modern game-theoretic algorithms. For simplicity, we often refer to a solution concept as the MSS that computes it when meaning is clear from the context. For example, we may say “Nash” to mean the MSS that computes NE. Moreover, the original version of PSRO specified deep RL as the best-response oracle [^1] PSRO actually works with any form of (approximate) response oracles, including search, planning, and evolutionary algorithms etc. For example, Li and Wellman Li and Wellman (2021) applied natural evolution strategies Wierstra et al. (2014) as the best response oracle for PSRO., enabling best-response computation in environments with a large number of states and actions. Finally, the payoffs of profiles in PSRO are estimated through simulation, and therefore a restricted game with estimated payoffs is usually called an empirical game. This idea of combining game-theoretic analysis with simulation in PSRO has been known as Empirical Game-Theoretic Analysis (EGTA) Wellman (2006), which provides a broad toolbox of techniques for game reasoning with game models based on simulation data. In an early EGTA work, Schvartzman and Wellman Schvartzman and Wellman (2009) deployed tabular RL as a best-response oracle (at a time when deep RL did not exist yet) and NE of the empirical game as a best-response target for strategy generation. PSRO extends tabular RL to deep RL and can be viewed as an instance of EGTA with strategies generated by iteratively best responding to MSS solutions using deep RL.
2.1 The Framework
A normal-form (aka strategic-form) representation of the full game $\mathcal{G}=(N,(S_{i}),(u_{i}))$ is a tuple, where $N$ is a finite set of players, each with a non-empty set of strategies $S_{i}$ and a utility function $u_{i}:\Pi_{j\in N}S_{j}\rightarrow\mathbb{R}$. A restricted game $\mathcal{G}_{S\downarrow X}$ is a projection of the full game $\mathcal{G}$, with players choosing from restricted strategy sets $X_{i}\subseteq S_{i}$. An empirical game $\hat{\mathcal{G}}_{S\downarrow X}=(N,(X_{i}),(\hat{u}_{i}))$ is a restricted game with simulated payoffs, where $\hat{u}_{i}$ represents the utility function estimated over the restricted strategy space $X$.
[…]
2.2 Strategy Exploration in PSRO
In essence, game-theoretic analysis in PSRO is performed by reasoning about empirical game models, induced from simulations run over combinations of strategies. To construct a feasible and effective game model for game analysis, the selection of the strategies is pivotal. In particular, a game model is expected to contain a much smaller number of strategies than the full game for representation tractability yet still maintain the key strategic information of the full game Balduzzi et al. (2018). This challenge of game model construction is described as the strategy exploration problem Jordan et al. (2010), which is the main research focus for developing PSRO methods. The goal of strategy exploration is to assemble an effective strategy portfolio for a game model with minimum computational cost (i.e., with the fewest strategies required). In PSRO, strategy exploration can be controlled by setting MSSs and ROs, which have a coupled impact; we refer to the joint choice as an MSS-RO combination.
[…]
In practice, the computation of regrets requires an exact best-response oracle, which is achievable in small games through methods such as strategy enumeration or dynamic programming. However, in large games, computing an exact best response becomes impractical. In such cases, approximate best responses are employed, providing a lower bound on regrets. The accuracy of regret estimation improves with a higher-quality oracle. With limited computational resources, when we cannot find a “better” response for any player, Oliehoek et al. Oliehoek et al. (2019) named the resulting profile a resource-bounded NE, with the interpretation: “with these resources, we did not refute that this is an NE”.
2.3 Organization of the Survey
As discussed in Section 2.2, a key design question for PSRO revolves around how to conduct effective strategy exploration given a specific goal. Although the performance of strategy exploration depends on the interplay between the chosen MSSs and ROs, existing literature predominantly focuses on setting either MSSs or ROs independently. Therefore, we organize our discussion on research directions and corresponding PSRO variants by first discussing setting MSSs and ROs independently (Sections 3 and 4 respectively). We then, in Section 5, discuss works that have investigated the joint choice of MSS-RO combination, before moving on to discuss in Section 6 how best to evaluate the effectiveness of such choices for strategy exploration. We then discuss research on improving the efficiency of PSRO (Section 7), explore applications of PSRO (Section 8), and conclude with open questions and future research directions (Section 9).
3 Strategy Exploration via MSS
In prior works, setting MSSs was a primary way to control strategy exploration. In this section, we discuss these works, their motivations, and their efficacy for strategy exploration.
3.1 PSRO with Normal-Form Empirical Games
In the standard PSRO framework, an empirical game is represented in normal form. While the simulations typically unfold through sequential observations and decisions over time, the empirical game abstracts away this temporal structure. This section deals with research questions and PSRO variants based on the normal-form representation.
3.1.1 Using Nash and its Variants as MSSs.
In PSRO, the most common MSS target is NE, which can be computed by various game-theoretic methods based on the normal-form empirical game. PSRO with NE is essentially DO with deep RL for computing (approximate) best responses. Therefore, PSRO with NE inherits the convergence property of DO, given mild assumptions about the quality of the best responses: In finite games, as long as beneficial deviations can always be found with non-zero probability, PSRO with NE as MSS target will converge to an NE in the limit.
The Overfitting Problem.
Despite its convergence guarantee in the limit, achieving exact convergence in large games is often unattainable due to constraints such as limited computational resources. Consequently, many prior works studying strategy exploration in large games revolves around the development of new algorithms that exhibit strong empirical performance (e.g., rapid convergence in terms of regret within a small number of PSRO iterations). The overfitting problem has two meanings for strategy exploration. First, it means that strategy exploration may overfit to the NE of the empirical game, which may not be an effective best-response target due to the limited information in the empirical game, potentially yielding ineffective strategy exploration. Second, it means overfitting to a specific equilibrium in general games (e.g., games with more than two players or general-sum games) without sufficiently exploring the whole strategy space. Note that the overfitting can be a problem for any solution concept; we discuss it in the context of NE specifically due NE’s prominence as the MSS target in the literature.
Regularization to Prevent Overfitting.
To address the overfitting problem, Lanctot et al. Lanctot et al. (2017) proposed an MSS, called Projected Replicator Dynamics (PRD), an adaptation of traditional replicator dynamics Taylor and Jonker (1978). PRD ensures a probability lower bound for selecting each strategy in the empirical game, allowing the new best response to train against not only strategies in the equilibrium support but also those outside the support. PRD can be viewed as a form of regularization to prevent overfitting the best response to the exact NE of the empirical game. Due to the diverse training targets, PRD also improves the stability of the new strategy.
Building on the concept of regularization, subsequent research has focused on designing MSSs that prevent overfitting by effectively regularizing the target profile. For instance, Wang et al. Wang et al. (2019) proposed an MSS that combines NE with a uniform distribution as the target profile, enabling the best response to an NE strategy mixed with exploration elements. Wright et al. Wright et al. (2019) developed a history-aware approach with a best-response (BR) target mixed with previous targets. Online double oracle Dinh et al. (2022) integrated PSRO with online learning to selected the BR target.
Wang and Wellman Wang and Wellman (2023b) addressed the overfitting problem by adopting an explicit view of regularization and introduced Regularized Replicator Dynamics (RRD), an MSS variant that truncates the NE search process in intermediate game models based on a regret criterion. Specifically, RRD computes the BR target by running RD in the empirical game, stopping once the regret of the current profile with respect to the empirical game meets a specified regret threshold. The regret criterion enables RRD to support direct control of the degree of regularization and can be adjusted to fit a specific game.
3.1.2 MSSs Beyond Nash
Rectified Nash
Balduzzi et al. Balduzzi et al. (2019) reformulated the strategy exploration problem as that of enlarging what they called the gamescape, which describes the strategy space covered by the empirical game. For zero-sum games, they proposed rectified Nash as an MSS designed to expand the gamescape and enhance a diversity measure called effective diversity. In rectified Nash, best responses are only applied to opponent’s equilibrium strategies that the learning player defeats or ties with.
Minimum-Regret Constrained Profile.
One notable observation for PSRO with NE is that the full-game regret, used as a measure for evaluating the performance of PSRO, of the empirical-game NE does not decrease monotonically over PSRO iterations. In the worst case, the full-game regrets will increase until the last iteration, when a full-game NE is found (one example can be found in the work by McAleer et al. McAleer et al. (2022b)). To address this issue, it was proposed to use minimum-regret constrained profiles (MRCP) Jordan et al. (2010); Wang et al. (2022) as an MSS. An MRCP is the profile with minimum regret with respect to the full game  [^2] MRCP was called the least-exploitable restricted distribution in the two-player zero-sum context by McAleer et al. McAleer et al. (2022b).. With MRCP as the MSS, the resulting PSRO variant is known as anytime PSRO because regret monotonically decreases as the empirical game grows. Despite the difficulty of computing MRCP in general games, anytime PSRO leverages the properties of two-player zero-sum games and computes MRCP by regret minimization against a best response (RM-BR) Johanson et al. (2012). In a further work McAleer et al. (2022a), anytime PSRO was extended by including not one but two strategies in the empirical game at each iteration, the first a best best response to MRCP and the other a best response to the other player’s latest strategy (i.e., the strategy added at the last PSRO iteration). This modification was observed to improve the performance of anytime PSRO. However, Wang et al. Wang et al. (2022) pointed out that MRCP regret will monotonically decrease for any MSS, which suggests that the changes to the MSS in anytime PSRO were not justified purely by the desire to monotonically decrease MRCP regret. We discuss the evaluation of strategy exploration further in Section 6.
Correlated Equilibrium.
Apart from the issues we discussed above, Marris et al. Marris et al. (2021) further argued that NE may not be an appropriate MSS for PSRO or even a solution concept in general-sum games due to its computational intractability. Therefore, they proposed utilizing correlated equilibrium (CE) and coarse correlated equilibrium (CCE) as MSSs, introducing a variant called Joint PSRO (JPSRO). Since multiple (C)CE exist in an empirical game, they select the unique (C)CE that maximizes the Gini impurity. Theoretical analysis demonstrated that JPSRO converges to a (C)CE. Relatedly, Team-PSRO finds TMECor equilibrium in two-team zero-sum games McAleer et al. (2023b).
MSSs Motivated by Specific Games.
In addition to the above well-established solution concepts, the literature on PSRO has also investigated other solution concepts, which originate from specific games but can be generally applied to various game settings. One example is the Risk-Averse Equilibrium (RAE) introduced by Slumbers et al. Slumbers et al. (2023), aimed at managing risk in multiagent systems. Specifically, RAE minimizes potential variance in rewards by accounting for the strategies of other players. Besides, Li et al. Li et al. (2023b) presented the Nash Bargaining Solution (NBS), a concept originating from the bargaining game settings, as an MSS. NBS can be computed by maximizing the product of player utilities in the empirical game.
Automated MSS Design.
Distinct from prior works that design MSSs based on various solution concepts and heuristics, Feng et al. Feng et al. (2021) proposed Neural Auto-Curricula (NAC) based on meta-learning, which automates the design of MSSs in an end-to-end manner. Specifically, MSSs in NAC are parametrized by a neural network, which is trained by minimizing the regret of the meta-strategy in the resulting empirical games. The resulting empirical games are generated through PSRO with the current MSS (i.e., the current neural network) in games sampled from a game distribution. With this training scheme, Feng et al. Feng et al. (2021) showed that NAC can learn an effective MSS for a family of games.
3.2 PSRO with Alternative Game Forms
Extensive-Form Games.
Instead of employing the normal-form representation, some PSRO variants use an extensive-form representation for the empirical game, offering a richer way to encompass temporal patterns in actions and information for underlying sequential games. One such example is given by extensive-form DO (XDO) McAleer et al. (2021), a DO variant with extensive-form game models. The extensive-form empirical game tree in XDO is constructed using the restricted set of players’ strategies. Similar to DO, NE is deployed as an MSS target, computed through Counterfactual Regret Minimization Zinkevich et al. (2007), and the best response computation will result in new actions at information states in the empirical game tree. Note that when a new action is added to an information state, multiple strategies will be added. So one iteration in XDO implicitly needs more simulation for profile evaluation than one DO iteration. Periodic DO Tang et al. (2023) extends XDO by improving the stopping threshold of the restricted game solver.
In a work by Konicki et al. Konicki et al. (2022), the benefits of leveraging the extensive-form representation for the empirical game were further explored. They showed that with an extensive-form representation in PSRO, the true game can be approximated more accurately than using a normal-form model constructed from the same amount of simulation data. This accuracy improvement stems from the fact that the simulation data for modeling a chance node in extensive form can be reused while the modeling needs to be re-simulated every time for evaluating a profile in normal form.
Mean-Field Games.
Muller et al. Muller et al. (2022) and Wang and Wellman Wang and Wellman (2023a) adapted PSRO to mean-field games (MFGs). Since the utility function for MFGs is not generally linear in the mean field, the empirical MFG model cannot be represented explicitly. Instead, Wang and Wellman Wang and Wellman (2023a) employed a game model learning approach Sokota et al. (2019), which is essentially a form of regression that learns a utility function over a restricted set of strategies and mean fields derived by these strategies. They proved the existence of NE in the empirical game with a restricted strategy set and the convergence of PSRO to NE in MFGs.
Diversity Measure Concept Diversity Type Enlargement Target Compatible MSS Behavioral Response Gamescape Policy Hull Nash $\alpha$-Rank Effective Diversity Balduzzi et al. (2019) Rectified Nash strategy Expected Cardinality Perez-Nieves et al. (2021) Determinantal point processes Convex Hull Enlargement Liu et al. (2021) Euclidean projection Occupancy Measure Mismatching Liu et al. (2021) f-divergence, occupancy measure Unified Diversity Measure Liu et al. (2022b) Strategy feature, diversity kernel Policy Space Diversity Yao et al. (2023) Bregman divergence, sequence-form
Table 1: Specifications of promoting-diversity methods.
4 Strategy Exploration via RO
Besides setting MSSs, establishing ROs to guide strategy exploration in PSRO has also been explored in the literature. One predominant way to design novel ROs is to include diversity measures in the standard RO, aiming at increasing the diversity of strategies in the empirical game.
For example, Perez-Nieves et al. Perez-Nieves et al. (2021) proposed diverse PSRO, which incorporates expected cardinality, a diversity measure defined through a determinantal point process, into the standard RO. Liu et al. Liu et al. (2021) define behavioral diversity (BD) and response diversity (RD), measuring diversity on different scales. BD is defined based on the distance between action-state coverages given by different strategies, while RD measures the distance between the payoff vector induced by the new strategy and the current empirical game. Subsequently, Liu et al. Liu et al. (2022b) proposed unified diversity measures to capture a variety of diversity metrics, which were later combined with the standard RO. Yao et al. Yao et al. (2023) observed that the current diversity measures for enlarging the gamescape fail to connect to the quality of NE approximation. They connected the RO to the quality of NE approximation by introducing population exploitability (PE), which is essentially the regret of MRCP Wang et al. (2022), to reflect the coverage of a policy hull (i.e., the convex combinations of strategies in the empirical game). They showed that a larger policy hull indicates lower PE. Therefore, they proposed an RO variant to enlarge the policy hull, aiming at promoting strategy exploration. We list the specifications of different methods for enhancing diversity in Table 1.
Aside from diversity, Li et al. Li et al. (2023b) deployed Monte Carlo tree search (MCTS) as the best-response oracle using different values (e.g., social welfare) to update values of nodes along the sample path in the back-propagation step of MCTS. The employment of different back-propagation values can also be viewed as modifications of the RO.
5 Strategy Exploration via Joint MSS-RO
Generally speaking, MSSs and ROs are not functionally interchangeable and can bring a coupling impact to strategy exploration, especially when a goal is specified. In this section, we discuss prior works that jointly vary MSSs and ROs.
$\alpha$-Rank.
Muller et al. Muller et al. (2020) proposed the adoption of $\alpha$-Rank Omidshafiei et al. (2019) as the preferred solution concept due to its computational scalability and uniqueness in many-player general-sum games. To make PSRO with $\alpha$-Rank as MSS converge, they introduced the preference-based best-response oracle, which essentially returns a set of strategies that maximizes the probability mass under $\alpha$-Rank from the set of better responses to the current strategy in a profile. With this MSS-RO combination, they proved that PSRO with $\alpha$-Rank converge to a concept, called sink strongly-connected components, which describes the distribution of strategies in long-term interactions.
Investigating the Joint Impact of MSSs and ROs.
An empirical study by Wang and Wellman Wang and Wellman (2024) explicitly investigated the coupling impact of MSSs and ROs in strategy exploration. This research experimented with a multitude of MSS and RO combinations with each RO possessing unique characteristics. Their experimental results underscore the pivotal role of ROs in steering strategy exploration towards desired objectives, such as higher social welfare. Moreover, they showed that with a careful selection of MSS, the performance of strategy exploration can be further improved.
6 Evaluating Strategy Exploration
In addition to designing novel strategy exploration algorithms, a significant effort has also been put into investigating methodological considerations in evaluating strategy exploration, and proposing and justifying new evaluation methods. In PSRO, it may seem natural to employ the same MSS for both strategy generation and evaluation, as much of the prior work in PSRO exploration has done. For example, if NE is used as the MSS for strategy generation, then the regret of NE of intermediate empirical games will be used as the performance measure at each iteration of PSRO. Similarly, the regret of a uniform distribution over strategies will be the performance measure when the uniform is employed as MSS, in which case PSRO reduces to fictitious play.
Wang and Wellman Wang et al. (2022) argued this evaluation approach could yield a misleading conclusion on the performance of MSS-RO combinations. They highlighted that each MSS-RO combination essentially generates a distinct sequence of strategies, and thus the empirical game model at any point reflects a distinct strategy space. The comparisons of different MSS-RO combinations are across different strategy spaces, which may not be faithfully represented by a simple summary such as an interim solution. Therefore, they proposed to use MRCP, the profile closest to the full-game NE (in regret) in the empirical game, as the evaluation metric for evaluating the performance of multiple MSS-RO combinations. This means that the MSS employed for strategy generation is independent of the MSS (e.g., MRCP) used for evaluation, which should be fixed when comparing different empirical games, regardless of the MSS with which they are generated. The requirement of fixing the evaluation MSS is named as the consistency criteria by the authors.
7 Improvements in Training Efficiency
There are two components in PSRO that are computationally demanding: best response computation (usually achieved by deep RL) and empirical game simulation. To improve the efficiency of PSRO, various methods have been developed, addressing issues related to these two aspects.
Parallelization.
Leveraging parallelization, Lanctot et al. Lanctot et al. (2017) proposed the Deep Cognitive Hierarchy (DCH) model, which creates a training hierarchy where each player trains a best response strategy (with deep RL) against the NE of the empirical game with strategies at the same level or below it. This warm-starts best response training and speeds up PSRO compared to training best responses from scratch. Motivated by DCH, McAleer et al. McAleer et al. (2020) proposed Pipeline PSRO (P2SRO). Similar to DCH, P2SRO initializes a bunch of strategies and assigns each strategy a level. Then P2SRO warm-starts training each strategy in parallel against the NE of the empirical game involving strategies with lower levels, which accelerates the overall training of PSRO. A major contribution of P2SRO is that it optimizes the training hierarchy, addressing issues with DCH such as heavy parallelization requirements and the need for an estimate of the number of PSRO iterations.
Sample Efficiency.
A distinctive characteristic of empirical game models is that they are derived or estimated from simulation data. To improve the sample efficiency of PSRO, Smith and Wellman Smith and Wellman (2023) proposed to learn a full-game model about the game dynamics from the simulator concurrently with running PSRO, aiming at reducing the simulation cost by querying the full-game model. Besides, Zhou et al. Zhou et al. (2022) developed an efficient PSRO (EPSRO) implementation for reducing the simulation cost of PSRO in two-player zero-sum games. The key insight is that the simulation for the empirical game is only used for computing best response target profiles. So as long as best response target profiles can be computed in other ways (e.g., a uniform MSS does not need the evaluation of a game model), there is no need to maintain the complete empirical game model, avoiding unnecessary simulations.
Transfer learning.
Transfer learning is a machine learning technique where a model trained on one task is repurposed for a different task. In PSRO, best-responding to different strategies can be viewed as such tasks and thus transfer learning can be applied to warm-start training new best responses. One example leveraging this idea is NeuPL Liu et al. (2022a), which represents all strategies in the strategy set via a shared neural network. NeuPL utilizes explicit parameter sharing for skill transfer, which was shown to effectively accelerate the adaptation to the opponent’s meta-strategy. Smith et al. Smith et al. (2023) also utilized transfer learning to reduce simulation costs in computing best responses. Their Mixed-Oracle method constructs a new best response by learning and maintaining best responses to the pure strategies of the opponent (represented as Q-value functions) and then mixing (Q-values) according to the meta-strategy.
8 Applications
Game-theoretic analysis in PSRO relies on game models, which abstract away the underlying game structures. This abstraction enables PSRO to solve a variety of games or address issues that can be formulated as a game, and yields numerous applications of PSRO in disparate domains. Specifically, PSRO has been applied to specialized games, including security games Wang et al. (2019); Wright et al. (2019); Fang (2019); Tong et al. (2020); Xu et al. (2021); Cui and Yang (2023), bargaining games Li et al. (2023b); Wang and Wellman (2024), Colonel Blotto games An and Zhou (2023), Pursuit-Evasion games Li et al. (2023a), auctions Li and Wellman (2021), and mechanism design Zhang et al. (2023).
Moreover, PSRO applications over time have been extended to real-world domains including anti-jamming in satellite communication Zou et al. (2022), decision-making in beyond-visual-range air combat Ma et al. (2019), solving the power imbalance in power system resilience Niu et al. (2021), and tackling the traveling salesman problem in combinatorial optimization Wang et al. (2021). Its utility is also evident in social network analysis for competitive influence maximization strategies Ansari et al. (2019) and in developing defense strategies for election safety Yin et al. (2018).
Besides the real-world applications, PSRO (including DO) has also been applied for designing novel algorithms in domains that can be modeled as a game. For example, the PSRO framework has facilitated developments in robust policy discovery within robust reinforcement learning domains Liang et al. (2023), value alignment in large language models Ma et al. (2023), public health services Killian et al. (2023), and the discovery of information in images Giboulot et al. (2023).
In the field of computer vision, several contributions towards data augmentation for object detection Behpour et al. (2019a), active learning Behpour et al. (2019b), semi-supervised multi-label classification Behpour (2018), and video tracking Fathony et al. (2018) underscore the PSRO framework’s impact on enhancing computer vision and structured prediction methodologies. Furthermore, PSRO-type methods have been adapted to enhance the training of Generative Adversarial Networks (GAN) Oliehoek et al. (2019); Aung et al. (2022), with notable implementations such as DO-GAN, which facilitates GAN training in high-dimensional spaces through strategy pruning and continual learning mechanisms.
9 Open Research Questions
The research directions outlined in the previous sections remain open-ended. Next, we describe several further research directions that we believe deserve the community’s attention.
Scalability in the Number of Players.
In the standard PSRO framework, an empirical game is represented in normal form. As the number of players increases, the normal-form representation expands exponentially, leading to a significant increase in the cost of evaluating the empirical game. While this problem can be mitigated in certain special game types such as symmetric games, the practicality of applying PSRO becomes impractical when dealing with a very large number of players. A potential remedy involves employing game model learning Sokota et al. (2019), essentially a utility function regression. The acquired utility function aims to extrapolate across the (empirical) strategy space, eliminating the need to evaluate each individual profile.
Another possible approach is to use polymatrix game representations. A polymatrix game is a many-player game where each player can be thought of as a node in an interaction graph and on every edge in this graph there is a bimatrix game that the node plays with its neighbours, gaining as utility the sum of payoffs across these bimatrix games. Polymatrix games have been used within equilibrium computation methods for many-player normal-form games, due to their linear formulation. In the work of both Govindan and Wilson Govindan and Wilson (2004) and Gemp at al. Gemp et al. (2022) a polymatrix approximation of the input game is maintained and updated (implicitly in the latter case). There appears to be potential to combine or extract elements from these methods to build polymatrix-based empirical game models for PSRO-type methods.
A Completely Parallel PSRO Implementation.
Both best-response computation and empirical game simulation can benefit from parallelization. Firstly, best-response computations among players are parallelizable as they are independent. Secondly, the evaluation of strategy profiles in the empirical game can be parallelized. Additionally, since the evaluation of each strategy profile involves running simulations multiple times and averaging the results, these simulation runs can also be parallelized. Consequently, the PSRO framework lends itself well to high parallelization. Despite this potential, there is currently no PSRO implementation that fully harnesses these three aspects.
Equilibrium Refinements.
As discussed in Section 3, PSRO can be adapted to calculating various solution concepts such as NE and CE. These solutions are often not unique, and the field of equilibrium refinement in game theory studies restrictions of these concepts to make the resulting solutions more realistic and better predictions of how the games will/should be played van Damme (2012). A natural research direction is to design PSRO variants to compute equilibrium refinements. For example, XDO, which uses an extensive-form representation, is a natural starting point to develop equilibrium refinement methods within extensive-form games (e.g. finding subgame-perfect equilibria). How can PSRO effectively compute these refinements by appropriately choosing the MSSs and ROs?
Multiple Equilibria.
Another important question relates to the existence and computation of multiple equilibria, particularly within general-sum games. While current PSRO research primarily focuses on identifying a sample equilibrium, the capability of PSRO to compute multiple equilibria remains under-explored. In principle, one could enumerate all equilibria in the restricted game, and use all in separate calls to the best-response oracles. In this way, we would be growing the restricted game in a way that encompasses all strategic information collected so far. This approach requires maintaining a larger restricted game, which inherently demands increased computational resources and necessitates more efficient computational frameworks to manage the enlarged strategy space; it is an interesting research question to decide how to most effectively split the use of limited computational resources between responding to multiple equilibria in a single PSRO iteration, versus being able to complete more of those iterations. Not all equilibria in the restricted game may be equally valuable for guiding strategy exploration. Can we predict this in advance?
Combining PSRO with Subgame Solving or CFR.
Other game-theoretic RL approaches, such as those based on counterfactual regret minimization (CFR) Brown et al. (2019); McAleer et al. (2023a) or policy gradients Fu et al. (2022); Perolat et al. (2022), could be potentially combined with PSRO. In contrast to PSRO, these approaches solve the entire game upfront, and as a result may be slower to converge on certain games and faster on others. CFR-based and policy-gradient-based methods might perform better in games where mixing carefully at many decision nodes is crucial, because PSRO usually mixes only at the root of the game and as a result might require many policies to effectively mix at information sets McAleer et al. (2021). Notably, XDO does mix at every decision node, and as a result is a great candidate for hybrid approaches with CFR-based and policy-gradient-based techniques. In particular, it would be useful to automatically decide on which decision nodes to mix. Since some games have sections that do not require mixing such as complex control tasks, those would be more efficient to use the deep RL policy from PSRO or XDO, while nodes that require careful mixing should defer to a CFR-based or policy-gradient-based technique.
Additionally, PSRO can be combined with subgame solving methods. Subgame solving methods, also called search methods, use extra compute at test time to analyze and improve a policy at a given information set without increasing exploitability. Some methods, such as depth-limited subgame solving Brown et al. (2018), already iteratively expand the strategy space by adding best responses. It would be interesting to combine the best responses learned in the full game via PSRO with those in the subgame that are used for subgame solving.
Automated Hyperparameter Tuning.
Hyperparameters exist universally in PSRO variants, including the lower bound of the probability of playing a strategy in PRD Lanctot et al. (2017), the regret threshold in RRD Wang and Wellman (2023b), the probability of employing the uniform MSS Wang et al. (2019), and the weights for diversity regularizers Perez-Nieves et al. (2021). These hyperparameters often require deliberate tuning to fit a specific game by running the same experiments multiple times for different hyperparameter choices, which can be costly. Although some of the hyperparameters can be set heuristically (e.g., by employing a simple annealing scheme), the question of how to tune the hyperparameters automatically and efficiently remains open. Specifically, can we design an automatic adjustment scheme for these hyperparameters such that PSRO can be directly applied to any game without manual game-specific fine turning?
PSRO with Large Language Models.
Recent achievements in large language models (LLMs) have drawn significant attention from various research communities. In the study of multiagent systems, researchers have explored diverse methods for integrating LLMs with game-theoretic principles, each serving distinct objectives. One research direction involves leveraging game-theoretic approaches to augment the capabilities of LLMs, such as enabling strategic responses in the presence of other agents Gemp et al. (2024), or fostering alignment with human values Ma et al. (2023). Conversely, LLMs can also be integrated into existing game-theoretic frameworks, expanding the range of applicable domains. Since PSRO provides a flexible game-theoretic framework, its potential to be combined with LLM in both ways should be noticed and investigated in future work.
Acknowledgement
We thank Michael P. Wellman for his invaluable advice on improving the accuracy and readability of the survey.
Call for Feedback
We encourage and welcome suggestions and comments from our readers. Despite our best efforts, there is always a possibility of oversight, including missed citations or overlooked information. Therefore, we invite our audience to engage with the content critically and to provide feedback if they notice any discrepancies or omissions. Your insights and contributions are invaluable in ensuring the accuracy and integrity of our work.